twice the square of a certain positive number added to 2 times the number make 84.find the number?
Answer by Cromlix(4381)

twice the square of a certain positive number added to 2 times the number make 84.
2 (x)^2 + 2(x) = 84
2x^2 + 2x = 84
2x^2 + 2x - 84 = 0
Divide thro'out by 2
x^2 + x - 42 + 0
(x + 7)(x - 6) = 0
x + 7 = 0
x = -7 No answer as negative
x - 6 = 0
x = 6 This is your answer.
2 (6)^2 + 2(6) = 84
2(36) + 12 = 84
72 + 12 = 84.
Hope this helps.
:-)